OpenStudy (anonymous):

What is the equation, in standard form, of a horizontal ellipse centered at (11, –6) with a major axis of length 16 and a minor axis of length 12

OpenStudy (anonymous):

\[\frac{ (x-11)^2 }{ 16 }+\frac{ (y+6)^2 }{ 12 }=1\]

OpenStudy (dumbcow):

no that is incorrect, the general form is: \[\frac{(x-11)^{2}}{a^{2}} + \frac{(y+6)^{2}}{b^{2}} = 1\] where major axis = 2a minor axis = 2b thus, a^2 = 64 and b^2 = 36